### v2

@@ A.1.2 Some technical lemmas
since change in any one of $\mathbf{z}_{\alpha}^{(v)}$ or $\mathbf{z}_{\alpha}^{(h)}$ affects $4$ finite difference terms in the summation. Combining the above controls, when $\mathbf{z}_{\alpha+}\in B_{1}(\mathbf{z}_{\alpha+}^{0}):=\{\mathbf{z}_{\alpha+}:\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}\leq 1\}$, it holds that
$\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})-$ $\log\mu(\mathbf{z}_{\alpha+}^{0}|\mathbf{z}_{\alpha-})+\mathbf{v}^{0}\cdot(\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0})$
$\geq$ $-\frac{L}{2}\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}^{2}-8\delta\left\|\mathbf{z}_{\alpha+}-\mathbf{z}_{\alpha+}^{0}\right\|_{2}\geq-\frac{L}{2}-8\delta.$
@@
$1=~{}$ $\int\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\mathrm{d}\mathbf{z}_{\alpha+}\geq\int_{B_{1}(\mathbf{z}_{\alpha+}^{0})}\exp\left(\log\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\right)\mathrm{d}\mathbf{z}_{\alpha+}$
@@
where we use the Jensen’s inequality and the symmetry of $B_{1}(\mathbf{z}_{\alpha-}^{0})$. Therefore,
$\mu(\mathbf{z}_{\alpha+}^{0}|\mathbf{z}_{\alpha-})\leq|B_{1}|^{-1}\exp\left(\frac{L}{2}+8\delta\right).$
This holds for arbitrary $\mathbf{z}_{\alpha+}^{0}$, so that the marginal distribution of $\mu(\mathbf{z}_{\alpha+})$ satisfies
$\mu(\mathbf{z}_{\alpha+})=\int\mu(\mathbf{z}_{\alpha+}|\mathbf{z}_{\alpha-})\mu(\mathbf{z}_{\alpha-})\mathrm{d}\mathbf{z}_{\alpha-}\leq|B_{1}|^{-1}\exp\left(\frac{L}{2}+8\delta\right)=:C^{\prime}.$
@@
$\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{-1}=~{}$ $\int_{\mathbb{R}^{2}}\frac{\mu(\mathbf{z}_{\alpha+})}{\left\|\mathbf{z}_{\alpha+}\right\|_{2}}\mathrm{d}\mathbf{z}_{\alpha}^{(v)}\mathrm{d}\mathbf{z}_{\alpha}^{(h)}\leq 1+\int_{\left\|\mathbf{z}_{\alpha+}\right\|_{2}\leq 1}\frac{C^{\prime}}{\left\|\mathbf{z}_{\alpha+}\right\|_{2}}\mathrm{d}\mathbf{z}_{\alpha}^{(v)}\mathrm{d}\mathbf{z}_{\alpha}^{(h)}$
@@
So that $C_{\pi}$ is dimension-independent. This completes the proof.
@@
In this proof, we work with the matrix representation $\mathbf{X}\in\mathbb{R}^{n\times n}$ for the unknown image such that the pixel at the index tuple $(\alpha,\beta)\in[n]\times[n]$ is given by $\mathbf{X}_{\alpha,\beta}$. Moreover, we use the notation $(\alpha\pm r)\coloneqq(\alpha-r,\alpha-r+1,\dots,\alpha+r-1,\alpha+r)$, and we let $w:(0\pm r)\times(0\pm r)\to\mathbb{R}_{>0}$ be the weights of the discrete PSF of the convolution kernel.
@@
(42) $\begin{split}V_{\alpha,\beta}(\mathbf{X})&=\frac{\lambda}{2}(\mathbf{Y}_{\alpha,\beta}-\sum_{\mu,\nu=-r}^{r}w(\mu,\nu)\mathbf{X}_{\alpha+\mu,\beta+\nu})^{2}\\ &~{}~{}~{}~{}+\delta\sqrt{(\mathbf{X}_{\alpha+1,\beta}-\mathbf{X}_{\alpha,\beta})^{2}+(\mathbf{X}_{\alpha,\beta+1}-\mathbf{X}_{\alpha,\beta})^{2}+\varepsilon}.\end{split}$
We now consider the full conditional of a fixed pixel $\pi(\mathbf{X}_{\hat{\alpha},\hat{\beta}}|\mathbf{X}_{\setminus(\hat{\alpha},\hat{\beta})})$, where $\setminus(\hat{\alpha},\hat{\beta})\coloneqq[n]\times[n]\setminus(\hat{\alpha},\hat{\beta})$. To find an expression for $\pi(\mathbf{X}_{\hat{\alpha},\hat{\beta}}|\mathbf{X}_{\setminus(\hat{\alpha},\hat{\beta})})$, it suffices to find all maximal cliques $V_{\alpha,\beta}$ which depend on $\mathbf{X}_{\hat{\alpha},\hat{\beta}}$. This follows from the Hammersley-Clifford theorem [11], and the fact that the posterior Eq. 14 is a Gibbs density. Therefore, let $\Theta_{\hat{\alpha},\hat{\beta}}\subseteq[n]\times[n]$ be the set of all maximal cliques $V_{\alpha,\beta}$ which depend on $\mathbf{X}_{\hat{\alpha},\hat{\beta}}$, and let $\Phi_{\hat{\alpha},\hat{\beta}}\subseteq[n]\times[n]$ be the set of pixels on which the cliques $\{V_{\alpha,\beta}\,|\,(\alpha,\beta)\in\Theta_{\hat{\alpha},\hat{\beta}}\}$ depend, i.e., $\Phi_{\hat{\alpha},\hat{\beta}}$ is the neighborhood of $\mathbf{X}_{\hat{\alpha},\hat{\beta}}$. Thus, we can write
$\pi(\mathbf{X}_{\hat{\alpha},\hat{\beta}}|\mathbf{X}_{\setminus(\hat{\alpha},\hat{\beta})})\propto\pi(\mathbf{X}_{\hat{\alpha},\hat{\beta}}|\mathbf{X}_{\Phi_{\hat{\alpha},\hat{\beta}}\setminus(\hat{\alpha},\hat{\beta})})\propto\exp{(-\sum_{(\alpha,\beta)\in\Theta_{\hat{\alpha},\hat{\beta}}}V_{\alpha,\beta}(\mathbf{X}_{\Phi_{\hat{\alpha},\hat{\beta}}}))},$
@@
In fact, $\Theta_{\hat{\alpha},\hat{\beta}}$ and $\Phi_{\hat{\alpha},\hat{\beta}}$ are determined by the first term in Eq. 42. The clique $V_{\alpha,\beta}$ depends on $\mathbf{X}_{\hat{\alpha},\hat{\beta}}$ if $\alpha=\hat{\alpha}-\mu$ and $\beta=\hat{\beta}-\nu$ where $\mu,\nu\in[-r,-r+1,\dots,r]$. Therefore, $\Theta_{\hat{\alpha},\hat{\beta}}=\hat{\alpha}\pm r\times\hat{\beta}\pm r$. From this follows, by checking the sum in the first term of Eq. 42, that $\Phi_{\hat{\alpha},\hat{\beta}}=\hat{\alpha}\pm 2r\times\hat{\beta}\pm 2r$.

### v3

@@ A.1.2 Some technical lemmas
since changing ${\bm{z}}_{\alpha}^{(v)}$ or ${\bm{z}}_{\alpha}^{(h)}$ affects $4$ finite difference terms in the summation. Combining the above controls, when ${\bm{z}}_{\alpha+}\in B_{1}({\bm{z}}_{\alpha+}^{0}):=\{{\bm{z}}_{\alpha+}:\left\|{\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0}\right\|_{2}\leq 1\}$, it holds that
$\log\mu({\bm{z}}_{\alpha+}|{\bm{z}}_{\alpha-})-$ $\log\mu({\bm{z}}_{\alpha+}^{0}|{\bm{z}}_{\alpha-})+{\bm{v}}^{0}\cdot({\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0})$
$\geq$ $-\frac{L}{2}\left\|{\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0}\right\|_{2}^{2}-8\delta\left\|{\bm{z}}_{\alpha+}-{\bm{z}}_{\alpha+}^{0}\right\|_{2}\geq-\frac{L}{2}-8\delta.$
@@
$1=~{}$ $\int\mu({\bm{z}}_{\alpha+}|{\bm{z}}_{\alpha-})\mathrm{d}{\bm{z}}_{\alpha+}\geq\int_{B_{1}({\bm{z}}_{\alpha+}^{0})}\exp\left(\log\mu({\bm{z}}_{\alpha+}|{\bm{z}}_{\alpha-})\right)\mathrm{d}{\bm{z}}_{\alpha+}$
@@
where we use the Jensen’s inequality and the symmetry of $B_{1}({\bm{z}}_{\alpha-}^{0})$. Therefore,
$\mu({\bm{z}}_{\alpha+}^{0}|{\bm{z}}_{\alpha-})\leq|B_{1}|^{-1}\exp\left(\frac{L}{2}+8\delta\right).$
This holds for arbitrary ${\bm{z}}_{\alpha+}^{0}$, so that the marginal distribution of $\mu({\bm{z}}_{\alpha+})$ satisfies
$\mu({\bm{z}}_{\alpha+})=\int\mu({\bm{z}}_{\alpha+}|{\bm{z}}_{\alpha-})\mu({\bm{z}}_{\alpha-})\mathrm{d}{\bm{z}}_{\alpha-}\leq|B_{1}|^{-1}\exp\left(\frac{L}{2}+8\delta\right)=:C^{\prime}.$
@@
$\mathbb{E}_{{\bm{x}}\sim\pi}\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{-1}=~{}$ $\int_{\mathbb{R}^{2}}\frac{\mu({\bm{z}}_{\alpha+})}{\left\|{\bm{z}}_{\alpha+}\right\|_{2}}\mathrm{d}{\bm{z}}_{\alpha}^{(v)}\mathrm{d}{\bm{z}}_{\alpha}^{(h)}\leq 1+\int_{\left\|{\bm{z}}_{\alpha+}\right\|_{2}\leq 1}\frac{C^{\prime}}{\left\|{\bm{z}}_{\alpha+}\right\|_{2}}\mathrm{d}{\bm{z}}_{\alpha}^{(v)}\mathrm{d}{\bm{z}}_{\alpha}^{(h)}$
@@
Thus, $C_{\pi}$ is dimension-independent. This completes the proof.
@@
In this proof, we work with the matrix representation ${\bm{X}}\in\mathbb{R}^{n\times n}$ for the unknown image such that the pixel at the index tuple $(\alpha,\beta)\in[n]\times[n]$ is given by ${\bm{X}}_{\alpha,\beta}$. Moreover, we use the notation $(\alpha\pm r)\coloneqq(\alpha-r,\alpha-r+1,\dots,\alpha+r-1,\alpha+r)$, and we let $w:(0\pm r)\times(0\pm r)\to\mathbb{R}_{>0}$ be the weights of the discrete PSF of the convolution kernel.
@@
(42) $\begin{split}V_{\alpha,\beta}({\bm{X}})&=\frac{\lambda}{2}({\bm{Y}}_{\alpha,\beta}-\sum_{\mu,\nu=-r}^{r}w(\mu,\nu){\bm{X}}_{\alpha+\mu,\beta+\nu})^{2}\\ &~{}~{}~{}~{}+\delta\sqrt{({\bm{X}}_{\alpha+1,\beta}-{\bm{X}}_{\alpha,\beta})^{2}+({\bm{X}}_{\alpha,\beta+1}-{\bm{X}}_{\alpha,\beta})^{2}+\varepsilon}.\end{split}$
We now consider the full conditional of a fixed pixel $\pi({\bm{X}}_{\hat{\alpha},\hat{\beta}}|{\bm{X}}_{\setminus(\hat{\alpha},\hat{\beta})})$, where $\setminus(\hat{\alpha},\hat{\beta})\coloneqq[n]\times[n]\setminus(\hat{\alpha},\hat{\beta})$. To find an expression for $\pi({\bm{X}}_{\hat{\alpha},\hat{\beta}}|{\bm{X}}_{\setminus(\hat{\alpha},\hat{\beta})})$, it suffices to find all maximal cliques $V_{\alpha,\beta}$ which depend on ${\bm{X}}_{\hat{\alpha},\hat{\beta}}$. This follows from the Hammersley-Clifford theorem [12] and the fact that the posterior Eq. 14 is a Gibbs density. Therefore, let $\Theta_{\hat{\alpha},\hat{\beta}}\subseteq[n]\times[n]$ be the set of all maximal cliques $V_{\alpha,\beta}$ which depend on ${\bm{X}}_{\hat{\alpha},\hat{\beta}}$, and let $\Phi_{\hat{\alpha},\hat{\beta}}\subseteq[n]\times[n]$ be the set of pixels on which the cliques $\{V_{\alpha,\beta}\,|\,(\alpha,\beta)\in\Theta_{\hat{\alpha},\hat{\beta}}\}$ depend, i.e., $\Phi_{\hat{\alpha},\hat{\beta}}$ is the neighborhood of ${\bm{X}}_{\hat{\alpha},\hat{\beta}}$. Thus, we can write
$\pi({\bm{X}}_{\hat{\alpha},\hat{\beta}}|{\bm{X}}_{\setminus(\hat{\alpha},\hat{\beta})})\propto\pi({\bm{X}}_{\hat{\alpha},\hat{\beta}}|{\bm{X}}_{\Phi_{\hat{\alpha},\hat{\beta}}\setminus(\hat{\alpha},\hat{\beta})})\propto\exp{(-\sum_{(\alpha,\beta)\in\Theta_{\hat{\alpha},\hat{\beta}}}V_{\alpha,\beta}({\bm{X}}_{\Phi_{\hat{\alpha},\hat{\beta}}}))},$
@@
In fact, $\Theta_{\hat{\alpha},\hat{\beta}}$ and $\Phi_{\hat{\alpha},\hat{\beta}}$ are determined by the first term in Eq. 42. The clique $V_{\alpha,\beta}$ depends on ${\bm{X}}_{\hat{\alpha},\hat{\beta}}$ if $\alpha=\hat{\alpha}-\mu$ and $\beta=\hat{\beta}-\nu$ where $\mu,\nu\in[-r,-r+1,\dots,r]$. Therefore, $\Theta_{\hat{\alpha},\hat{\beta}}=\hat{\alpha}\pm r\times\hat{\beta}\pm r$. From this follows, by checking the sum in the first term of Eq. 42, that $\Phi_{\hat{\alpha},\hat{\beta}}=\hat{\alpha}\pm 2r\times\hat{\beta}\pm 2r$.